OpenStudy (anonymous):

Identify the Exponent Law shown in the example below. (ab)3 = a3b3 Power of a Product Product of Powers Power of a Power None of the Above

jimthompson5910 (jim_thompson5910):

(...)^3 is a power a*b is a product

jimthompson5910 (jim_thompson5910):

combine the two, in the right order, to say that (a*b)^3 is a power of a product

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

can you help me with more or do I have to close this one and put another question up?

jimthompson5910 (jim_thompson5910):

I don't mind either way

jimthompson5910 (jim_thompson5910):

do what works best for you

OpenStudy (anonymous):

(5x4y2z2)(3x4y3z5). 15x16y6z10 15x8y5z7 8x8y5z7 8x16y6z10 wouldn't it be c?

jimthompson5910 (jim_thompson5910):

5*3 isn't 8, it's 15

jimthompson5910 (jim_thompson5910):

only the exponents add, the coefficients multiply

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

No its A

jimthompson5910 (jim_thompson5910):

exponent for x: 4+4 = 8 exponent for y: 2+3 = 5 exponent for z: 2+5 = 7

jimthompson5910 (jim_thompson5910):

hmm decisions decisions...A or B

OpenStudy (anonymous):

its B

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

do you do the same thing for this? Choose the correct simplification of the expression (2ab5)3. 8ab15 8a3b8 5a3b15 8a3b15

jimthompson5910 (jim_thompson5910):

you would use the rule given at the top of the page

jimthompson5910 (jim_thompson5910):

(a*b)^c = a^c * b^c

jimthompson5910 (jim_thompson5910):

so this means that (2ab^5)^3 = 2^3 * a^3 * (b^5)^3 = ???

OpenStudy (anonymous):

so D...?

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

you're a pro at this already

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

im not sure about this one Choose the correct simplification of the expression (3xy4)2(y2)3. 6x2y14 9x2y14 9x3y11 6x3y11

jimthompson5910 (jim_thompson5910):

simplify (3xy^4)^2 first and tell me what you get

OpenStudy (anonymous):

7xy8?

jimthompson5910 (jim_thompson5910):

hmm no unfortunately

jimthompson5910 (jim_thompson5910):

(3xy^4)^2 = 3^2 * x^2 * (y^4)^2 (3xy^4)^2 = 9 * x^2 * y^8

jimthompson5910 (jim_thompson5910):

what do you get when you simplify (y^2)^3

OpenStudy (anonymous):

y16

jimthompson5910 (jim_thompson5910):

y^6 so you are close

jimthompson5910 (jim_thompson5910):

(3xy^4)^2(y2)^3 becomes 9x^2*y^8*y^6 now simplify 9x^2*y^8*y^6

OpenStudy (anonymous):

18xy16y12.. I'm doing this wrong.. :/

jimthompson5910 (jim_thompson5910):

all you do at this point is add the y exponents

jimthompson5910 (jim_thompson5910):

8+6 = 14

OpenStudy (anonymous):

so 18xy14

jimthompson5910 (jim_thompson5910):

so the final answer is \[\large 9x^{2}y^{14}\]

jimthompson5910 (jim_thompson5910):

idk how you're getting 18

OpenStudy (anonymous):

idk either lol

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

well that 2 is an exponent and you're not multiplying it by 9 to get 18

OpenStudy (anonymous):

so what do you do with it?

jimthompson5910 (jim_thompson5910):

you just leave the 9 as it is

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

what about this? is it D or B? Choose the correct simplification of the expression (−13g2h5k2)0. 1 0 13g2h5k2 −13

jimthompson5910 (jim_thompson5910):

0 is the exponent right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

anything to the 0th power (except 0 itself) is always 1

jimthompson5910 (jim_thompson5910):

ex: 2^0 = 1 9^0 = 1 115^0 = 1 you get the idea

jimthompson5910 (jim_thompson5910):

so if x is not zero, then x^0 = 1

OpenStudy (anonymous):

ohh okay

jimthompson5910 (jim_thompson5910):

so you could have some really long and really ridiculous expression raised to the 0th power....and it would all boil down to 1

OpenStudy (anonymous):

I don't feel like doing that lol

jimthompson5910 (jim_thompson5910):

yeah just know that (−13g^2h^5k^2)^0 = 1 and you're good to go

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw glad to be of help